OpenStudy (anonymous):

20% acid solution is mixed with a 70% acid solution to get 50 liters of a 40% solution. How much of the 70% solution is used?

hero (hero):

Hint: x + y = 50 .20x + .70y = 20

hero (hero):

Solve the system and find y

hero (hero):

Chero, you there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry. I am doing something wrong because I am getting y=28.57-20x/.70

hero (hero):

1. Isolate x in each equation x = 50 - y x = (20 - .70y)/.20 1. Set x = x x = x 50 - y = (20 - .70y)/.20 2. Cross multiply: .20(50 - y) = 20 - .70y 3. Simplify and solve for y: 10 - .20y = 20 - .70y .70y - .20y = 20 - 10 .50y = 10 y = 10/.50 y = 20 Since y represents the 70% solution, 20 liters of that solution was used.

OpenStudy (anonymous):

THank you soooo much! I've being doing math since 10 am this morning! Tomorrow is our last day to finish this stuff so I really appreciate your help!!! I am just about fried!

hero (hero):

You've been doing math since 10am for your son?